OpenStudy (babe2323433):

What is the simplified form of the quantity y−squared minus y minus 12 over the quantity y−squared plus 8y plus 15?

OpenStudy (babe2323433):

@sunnnystrong

OpenStudy (sunnnystrong):

hmmm we just did this one XD ^^

OpenStudy (babe2323433):

Nooo we did one similar but not this one..

OpenStudy (babe2323433):

@sunnnystrong

OpenStudy (babe2323433):

This is the last question

OpenStudy (sunnnystrong):

\[\frac{ y^2 -y-12}{ y^2+8y+15 }=\frac{ (y-4)(y+3) }{ (y+3)(y+5) }\]

OpenStudy (sunnnystrong):

hmm.. lookks pretty familiar lol

OpenStudy (sunnnystrong):

but basically what's next? @babe2323433 what cancels? what's the answer?

OpenStudy (babe2323433):

(y+3) cancels and the answer is (y-4)/(y+5)

OpenStudy (sunnnystrong):

yep :D good job!